# Kummer-Artin-Schreier-Witt Theory

### v2

@@ 5.2 When $s>1$
Here we set $\mathbb{P}^{s}_{R}=\operatorname{Proj}R[X_{0},X_{1},\ldots,X_{s-1},V_{0}]$ and
@@
where the polynomials $G_{i-1}(X_{0},\ldots,X_{i-2})$ are computed in Section 3, and $L_{i+1}/L_{i}$ is given by
@@
where $S_{i-1}$ is an Artin-Schreier-Witt type polynomial similar to (31) in the ring
@@
for $i=1,\ldots,s-1$. Assume the induction hypothesis that $\mathcal{W}_{s-1}=\operatorname{Spec}B_{s-1}$, with $B_{s-1}$ given by
$R\left[Y_{0},\ldots,Y_{s-2},X_{s-1},\frac{1}{1+\lambda Y_{0}},\ldots,\frac{1}{E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2}},\frac{1}{G_{s}(S_{0},\ldots,S_{i-2})+p\lambda X_{s-1}}\right],$
and the $S_{j}(Y_{0},\ldots,Y_{j})$ as defined above.
@@
where $E_{s}(Y_{0},\ldots,Y_{s-2})$ is defined in Section 4, transforming equation (32) when $i=s-1$ into
$(E_{s}(Y_{0},\ldots,Y_{s-2})+\lambda Y_{s-1})^{p}=(E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2})\left(G_{s}(S_{0},\ldots,S_{i-2})+p\lambda X_{s-1}\right),$ (33)
@@
$\left(\frac{1}{E_{s}(Y_{0},\ldots,Y_{s-2})+\lambda Y_{s-1}}\right)^{p}=\frac{1}{G_{s}(S_{0},\ldots,S_{i-2})+p\lambda X_{s-1}}\cdot\frac{1}{E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2}}.$
@@
These equations imply that $Y_{s-1}$ and $\frac{1}{E_{s}+\lambda Y_{s-1}}$ are integral over $B_{s-1}$, and that $X_{s-1}$ lies in $B_{s}$. Therefore, by reasoning similar to that in the proof of Proposition 5.1, we conclude that the integral closure of $B_{s-1}$ in $L_{s}$ is $B_{s}$. ∎
Thus, $B_{s}$ is also the integral closure of $\mathcal{O}_{\mathcal{V}_{s}}$ in $L_{s}$, and $\operatorname{Spec}B_{s}=:\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}$ is étale with a Galois group $\mathbb{Z}/p^{s}$ and has good reduction. In particular, the special fiber $\operatorname{Spec}\mathbb{F}_{p}[y_{0},\ldots,y_{s-1}]\rightarrow\operatorname{Spec}\mathbb{F}_{p}[x_{0},\ldots,x_{s-1}]$ corresponds to the Witt vector $(x_{0},\ldots,x_{s-1})$.
@@ 6 The Group Structure
In this section, set $R:=\mathbb{Z}_{(p)}[\zeta_{p^{s}}]$, $\hat{R}:=\mathbb{Z}_{(p)}\llbracket\zeta_{p^{s}}-1\rrbracket$, and $\pi:=\zeta_{p^{s}}-1$. From Proposition 4.1 and Proposition 3.2, we have the (not necessarily algebraic) coordinate maps $\gamma:\mathcal{W}_{s}(R)\rightarrow W_{s}(\hat{R})$ and $\tau:\mathcal{V}_{s}(R)\rightarrow W_{s}(\hat{R})$, given by
$(Y_{0},\ldots,Y_{s-1})\mapsto(Y_{0}^{\prime},\ldots,Y_{s-1}^{\prime})$ (34)
@@
$(X_{0},\ldots,X_{s-1})\mapsto(X_{0}^{\prime},\ldots,X_{s-1}^{\prime}),$ (35)
respectively. More precisely, each $X_{i}^{\prime}$ is computed inductively as follows:
@@
Conversely, given $(Y_{0}^{\prime},\ldots,Y_{s-1}^{\prime})\in W_{s}(\hat{R})$ (resp. $(X_{0}^{\prime},\ldots,X_{s-1}^{\prime})\in W_{s}(\hat{R})$), one can compute the corresponding point in $\mathcal{W}_{s}(\hat{R})$ (resp. $\mathcal{V}_{s}(\hat{R})$) as follows:
$Y_{0}$ $=\frac{1}{\lambda}\left(\exp\left(\lambda Y_{0}^{\prime}\right)-1\right),$
@@
). We can therefore naturally declare the group operations on $\mathcal{W}_{s}(\hat{R})$ and $\mathcal{V}_{s}(\hat{R})$.
@@
Define a group structure on $\mathcal{W}_{s}$ (and similarly on $\mathcal{V}_{s}$) over $\hat{R}$ by pulling back the Witt vector group structure on $W_{s}$ via the map $\gamma$ (or via $\tau$).
Remark 6.2.
Specifically, suppose $(X_{0,1},\ldots,X_{s-1,1})$ and $(X_{0,2},\ldots,X_{s-1,2})$ are points in $\mathcal{V}_{s}(\hat{R})$ with associated Witt vectors $(X_{0,1}^{\prime},\ldots,X_{s-1,1}^{\prime})$ and $(X_{0,2}^{\prime},\ldots,X_{s-1,2}^{\prime})$, respectively. Then we define their sum $(X_{0,3},\ldots,X_{s-1,3})$ as the unique point in $\mathcal{V}_{s}(\hat{R})$ satisfying
$\tau(X_{0,3},\ldots,X_{s-1,3})=(X_{0,1}^{\prime},\ldots,X_{s-1,1}^{\prime})+(X_{0,2}^{\prime},\ldots,X_{s-1,2}^{\prime}),$
where the right-hand side denotes the sum of Witt vectors in $W_{s}$.
@@
Definition 6.1 also makes $\psi_{s}:\mathcal{W}_{s}(R)\to\mathcal{V}_{s}(R)$ an isogeny.

### v3

@@ 5.2 The case $s>1$
Here we set $\mathbb{P}^{s}_{R}=\operatorname{Proj}R[X_{0},X_{1},\ldots,X_{s-1},Y]$ and
@@
where the polynomials $G_{2},\ldots,G_{s}$ are computed in Section 3, and $L_{i+1}/L_{i}$ is given by
@@
where $S_{i-1}$ is an Artin–Schreier–Witt type polynomial similar to (34). For $i\geq 2$, it lies in the ring
@@
for $i=2,\ldots,s-1$. Following the same process as in Section 5.1, one shows that $\mathcal{W}_{1}=\operatorname{Spec}B_{1}$, where $B_{1}$ is given by
$R\left[Y_{0},X_{1},\ldots,X_{s-1},\frac{1}{1+\lambda Y_{0}},\frac{1}{G_{2}(S_{0})+p\lambda X_{1}},\ldots,\frac{1}{G_{s}(S_{0},\ldots,X_{s-2})+p\lambda X_{s-1}}\right]$
Here $\mathcal{W}_{1}$ denotes the first stage in the tower over $\mathcal{V}_{s}$ (i.e. the pullback of the $s=1$ construction along the natural projection $\mathcal{V}_{s}\to\mathcal{V}_{1}$).
Assume the induction hypothesis that $\mathcal{W}_{s-1}=\operatorname{Spec}B_{s-1}$, with $B_{s-1}$ given by
$R\left[Y_{0},\ldots,Y_{s-2},X_{s-1},\frac{1}{1+\lambda Y_{0}},\ldots,\frac{1}{E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2}},\frac{1}{G_{s}(S_{0},\ldots,S_{s-2})+p\lambda X_{s-1}}\right],$
and the polynomials $S_{j}(Y_{0},\ldots,Y_{j})$ are as defined above.
@@
where $E_{s}(Y_{0},\ldots,Y_{s-2})$ is defined in Section 4, transforming equation (35) when i=s1i=s-1 into
$(E_{s}(Y_{0},\ldots,Y_{s-2})+\lambda Y_{s-1})^{p}=(E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2})\left(G_{s}(S_{0},\ldots,S_{s-2})+p\lambda X_{s-1}\right),$ (36)
@@
$\left(\frac{1}{E_{s}(Y_{0},\ldots,Y_{s-2})+\lambda Y_{s-1}}\right)^{p}=\frac{1}{G_{s}(S_{0},\ldots,S_{s-2})+p\lambda X_{s-1}}\cdot\frac{1}{E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2}}.$
@@
These equations imply that $Y_{s-1}$ and $\frac{1}{E_{s}+\lambda Y_{s-1}}$ are integral over $B_{s-1}$, and that $X_{s-1}$ lies in $B_{s}$. Hence $B_{s}$ is integral over $B_{s-1}$ and contained in the integral closure $\widetilde{B_{s-1}}$ of $B_{s-1}$ in $L_{s}$. Conversely, as in Proposition 5.1, $B_{s}$ is normal (being a localization of a polynomial ring over the DVR $R$), contains $B_{s-1}$, and has fraction field $L_{s}$; therefore $\widetilde{B_{s-1}}\subseteq B_{s}$. Thus $\widetilde{B_{s-1}}=B_{s}$. ∎
Corollary 5.3.
The special fiber of $\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}$ is isomorphic to the morphism $\wp$ in (8). In particular, it is unramified.
Proof.
By Proposition 5.2, $B_{s}$ is also the integral closure of $\mathcal{O}_{\mathcal{V}_{s}}$ in $L_{s}$. Furthermore, the special fiber of $\mathcal{W}_{s}$ (resp. $\mathcal{V}_{s}$) is $\mathbb{F}_{p}[y_{0},\ldots,y_{s-1}]$ (resp. $\mathbb{F}_{p}[x_{0},\ldots,x_{s-1}]$). Together with Corollary 4.13, this shows that the special fiber is isomorphic to $\wp$ in (8). Finally, $\wp$ is unramified (equivalently étale) on $W_{s}$ since $\ker(\wp)\cong\mathbb{Z}/p^{s}$ is finite étale by (8). ∎
Proposition 5.4.
The cover $\mathcal{W}_{s}\to\mathcal{V}_{s}$ is an étale cover with good reduction.
Proof.
The assertion of good reduction holds because the special fiber is an étale $\mathbb{Z}/p^{s}$-torsor between smooth $k$-schemes, as established in Corollary 5.3.
For flatness, fix $x\in\mathcal{W}_{s}$ mapping to $y\in\mathcal{V}_{s}$, and set $A:=\mathcal{O}_{\mathcal{V}_{s},y}$ and $B:=\mathcal{O}_{\mathcal{W}_{s},x}$. Since $\mathcal{V}_{s}$ and $\mathcal{W}_{s}$ are smooth over the DVR $R$, both $A$ and $B$ are regular local rings, hence $B$ is Cohen–Macaulay. The morphism $\mathcal{W}_{s}\to\mathcal{V}_{s}$ is finite, so $A\to B$ is integral, and in particular $\dim A=\dim B$. By Miracle Flatness (Stacks Project, Lemma 10.128.1), it follows that $B$ is flat over $A$. Hence $\mathcal{W}_{s}\to\mathcal{V}_{s}$ is finite and flat.
For étaleness, let $M:=\Omega_{\mathcal{W}_{s}/\mathcal{V}_{s}}$. Since $\mathcal{W}_{s}\to\mathcal{V}_{s}$ is of finite type, $M$ is a coherent $\mathcal{O}_{\mathcal{W}_{s}}$-module; since the morphism is finite, $M$ is in fact finite over $\mathcal{O}_{\mathcal{W}_{s}}$. By Proposition 3.11, the generic fiber is unramified, so $M[1/\pi]=0$. Thus, for each $x$, the stalk $M_{x}$ is a finitely generated $B$-module with $(M_{x})[1/\pi]=0$, hence there exists $n\geq 1$ such that $\pi^{n}M_{x}=0$. By Corollary 5.3, the special fiber is unramified, so $M_{x}/\pi M_{x}=0$, i.e. $M_{x}=\pi M_{x}$. Therefore $M_{x}=\pi^{n}M_{x}=0$ for all $x$, so $M=0$ and the morphism is unramified. Since it is also flat, it is étale. ∎
@@ 6 The Group Structure
In this section, we explain how the covers constructed in the previous sections carry natural group scheme structures, and how the morphism $\psi_{s}\colon\mathcal{W}_{s}\to\mathcal{V}_{s}$ becomes a group homomorphism. The key point is that, after completing along the identity sections, certain explicit coordinate maps (defined below) identify the formal neighborhoods of $\mathcal{V}_{s}$ and $\mathcal{W}_{s}$ with the formal Witt vector scheme; we then transport Witt vector addition to obtain formal group laws, and finally descend these formal structures to algebraic Hopf algebras over $R$.
Let $R:=\mathbb{Z}_{(p)}[\zeta_{p^{s}}]$ be a discrete valuation ring, and let $\pi:=\zeta_{p^{s}}-1$ be its uniformizer.
Coordinate rings. Let
$A:=\mathcal{O}(\mathcal{V}_{s})=R[X_{0},\dots,X_{s-1}]\bigl[U_{0}^{-1},\dots,U_{s-1}^{-1}\bigr],$
with defining relations
$U_{0}=1+p\lambda X_{0},\qquad U_{j}=G_{j+1}(X_{0},\dots,X_{j-1})+p\lambda X_{j}\quad(1\leq j\leq s-1),$
@@
$B:=\mathcal{O}(\mathcal{W}_{s})=R[Y_{0},\dots,Y_{s-1}]\bigl[Z_{0}^{-1},\dots,Z_{s-1}^{-1}\bigr],$
with defining relations
$Z_{0}=1+\lambda Y_{0},\qquad Z_{j}=E_{j+1}(Y_{0},\dots,Y_{j-1})+\lambda Y_{j}\quad(1\leq j\leq s-1).$
[…]
$\widehat{A}\cong R\llbracket X_{0},\ldots,X_{s-1}\rrbracket,\qquad\widehat{B}\cong R\llbracket Y_{0},\ldots,Y_{s-1}\rrbracket.$
Let $e_{\mathcal{V}}\in\mathcal{V}_{s}(R)$ and $e_{\mathcal{W}}\in\mathcal{W}_{s}(R)$ denote the identity sections. The formal completions of $\mathcal{V}_{s}$ and $\mathcal{W}_{s}$ along these sections are
[…]
Similarly, letting $0\in W_{s}(R)$ be the identity section, we write
$\widehat{W}_{s}:=\widehat{W}_{s,0}=\operatorname{Spf}\bigl(\widehat{\mathcal{O}}_{W_{s},0}\bigr)$
for the formal completion of $W_{s}$ at 0.
From Proposition 4.1 and Proposition 3.4, we have the (not necessarily algebraic) coordinate maps $\gamma$ and $\tau$. On formal completions along the identity sections, $\tau$ maps $\widehat{\mathcal{V}}_{s}$ to $\widehat{W}_{s}$, and $\gamma$ maps $\widehat{\mathcal{W}}_{s}$ to $\widehat{W}_{s}$. Concretely,
$\gamma\colon\widehat{\mathcal{W}}_{s}\longrightarrow\widehat{W}_{s},\qquad(Y_{0},\ldots,Y_{s-1})\longmapsto(Y_{0}^{\prime},\ldots,Y_{s-1}^{\prime})$ (37)
$\tau\colon\widehat{\mathcal{V}}_{s}\longrightarrow\widehat{W}_{s},\qquad(X_{0},\ldots,X_{s-1})\longmapsto(X_{0}^{\prime},\ldots,X_{s-1}^{\prime}).$ (38)
More precisely, each $X_{i}^{\prime}$ is computed inductively as follows:
@@
Conversely, given $(Y_{0}^{\prime},\ldots,Y_{s-1}^{\prime})\in\widehat{W}_{s}$ (resp. $(X_{0}^{\prime},\ldots,X_{s-1}^{\prime})\in\widehat{W}_{s}$), one can compute the corresponding point in the formal completion $\widehat{\mathcal{W}}_{s}$ (resp. $\widehat{\mathcal{V}}_{s}$) along the identity section as follows.
$Y_{0}$ $=\frac{1}{\lambda}\left(\exp\left(\lambda Y_{0}^{\prime}\right)-1\right),$ (39)
@@
). We denote the induced identification $\widehat{W}_{s}\to\widehat{\mathcal{V}}_{s}$ by $\tau^{-1}$ (resp. the identification $\widehat{W}_{s}\to\widehat{\mathcal{W}}_{s}$ by $\gamma^{-1}$).
Lemma 6.1.
With notation as above, we have the following.
  1. (a)
    The coordinate map $\tau$ defined in (38) restricts to an isomorphism of formal schemes over $R$, which we denote
    $\widehat{\tau}\colon\widehat{\mathcal{V}}_{s}\xrightarrow{\ \sim\ }\widehat{W}_{s},$
    whose inverse is induced by the $X$-formulas in (40). Equivalently, on completed local rings one gets mutually inverse continuous $R$-algebra homomorphisms
    $\widehat{\tau}^{\sharp}\colon\widehat{\mathcal{O}}_{W_{s},0}\longrightarrow\widehat{\mathcal{O}}_{\mathcal{V}_{s},e_{\mathcal{V}}},\qquad(\widehat{\tau}^{-1})^{\sharp}\colon\widehat{\mathcal{O}}_{\mathcal{V}_{s},e_{\mathcal{V}}}\longrightarrow\widehat{\mathcal{O}}_{W_{s},0}.$
  2. (b)
    The coordinate map $\gamma$ defined in (37) restricts to an isomorphism of formal schemes over $R$, which we denote
    $\widehat{\gamma}\colon\widehat{\mathcal{W}}_{s}\xrightarrow{\ \sim\ }\widehat{W}_{s},$
    whose inverse is induced by the $Y$-formulas in (39). Equivalently, on completed local rings one gets mutually inverse continuous $R$-algebra homomorphisms
    $\widehat{\gamma}^{\sharp}\colon\widehat{\mathcal{O}}_{W_{s},0}\longrightarrow\widehat{\mathcal{O}}_{\mathcal{W}_{s},e_{\mathcal{W}}},\qquad(\widehat{\gamma}^{-1})^{\sharp}\colon\widehat{\mathcal{O}}_{\mathcal{W}_{s},e_{\mathcal{W}}}\longrightarrow\widehat{\mathcal{O}}_{W_{s},0}.$
Proof.
We prove (a); the proof of (b) is identical, replacing $p\lambda$ with $\lambda$ and $G_{j+1},G_{j+1,p}$ with $E_{j+1},E_{j+1,p}$.
Let $I_{\mathcal{V}}:=(X_{0},\ldots,X_{s-1})\subset A$ and $I_{W}:=(X_{0}^{\prime},\ldots,X_{s-1}^{\prime})\subset\mathcal{O}(W_{s})$ be the ideals of the identity sections. By the discussion above, the $I_{\mathcal{V}}$-adic completion of $A$ identifies with the completed local ring of $\mathcal{V}_{s}$ at $e_{\mathcal{V}}$, i.e.
$\widehat{A}\cong\widehat{\mathcal{O}}_{\mathcal{V}_{s},e_{\mathcal{V}}}\cong R\llbracket X_{0},\ldots,X_{s-1}\rrbracket,\qquad\widehat{\mathcal{O}}_{W_{s},0}\cong R\llbracket X_{0}^{\prime},\ldots,X_{s-1}^{\prime}\rrbracket.$
Step 1: $\tau$ defines a morphism of formal schemes. To define a morphism $\widehat{\tau}:\widehat{\mathcal{V}}_{s}\to\widehat{W}_{s}$, it suffices to give a continuous $R$-algebra map
$\widehat{\tau}^{\sharp}:R\llbracket X_{0}^{\prime},\ldots,X_{s-1}^{\prime}\rrbracket\to R\llbracket X_{0},\ldots,X_{s-1}\rrbracket$
sending $I_{W}$ into $I_{\mathcal{V}}$.
For $j=0$ set
$X_{0}^{\prime}=\frac{1}{p\lambda}\ln(1+p\lambda X_{0})=\sum_{n\geq 1}(-1)^{n-1}\frac{(p\lambda)^{n-1}}{n}X_{0}^{n}.$
Since $p\lambda\in pR$, one has $\nu((p\lambda)^{n-1})\geq n-1\geq\nu(n)$, so the coefficients lie in $R$ and $X_{0}^{\prime}\in I_{\mathcal{V}}$.
Assume inductively that $X_{0}^{\prime},\ldots,X_{j-1}^{\prime}\in R\llbracket X_{0},\ldots,X_{j-1}\rrbracket$ and have zero constant term. Define
$X_{j}^{\prime}=\frac{1}{p\lambda}\ln\!\Bigl(\frac{G_{j+1}(X_{0},\ldots,X_{j-1})+p\lambda X_{j}}{G_{j+1,p}(X_{0}^{\prime},\ldots,X_{j-1}^{\prime},0)}\Bigr).$
By Proposition 3.8,
$G_{j+1}(X_{0},\ldots,X_{j-1})\equiv G_{j+1,p}(X_{0}^{\prime},\ldots,X_{j-1}^{\prime},0)\pmod{p\lambda}$.
Indeed, $G_{j+1,p}(X_{0}^{\prime},\ldots,X_{j-1}^{\prime},0)-G_{j+1}(X_{0},\ldots,X_{j-1})=G^{\prime}_{j+1}$ has $\nu(G^{\prime}_{j+1})>\frac{p}{p-1}=\nu(p\lambda)$ by Proposition 3.8, hence $G^{\prime}_{j+1}\in(p\lambda)$. Moreover $G_{j+1,p}(X_{0}^{\prime},\ldots,X_{j-1}^{\prime},0)$ has constant term $1$, hence is a unit in the complete local ring $R\llbracket X_{0},\ldots,X_{j-1}\rrbracket$. Therefore the ratio is of the form $1+p\lambda\Delta$ with $\Delta\in R\llbracket X_{0},\ldots,X_{j}\rrbracket$, and since numerator and denominator both have constant term $1$, one has $\Delta\in I_{\mathcal{V}}\,R\llbracket X_{0},\ldots,X_{j}\rrbracket$.
Thus $\ln(1+p\lambda\Delta)$ is a well-defined formal power series in the $I_{\mathcal{V}}$-adic topology, and dividing by $p\lambda$ yields $X_{j}^{\prime}\in I_{\mathcal{V}}$. This defines a continuous $R$-algebra map $\widehat{\tau}^{\sharp}$.
Step 2: $\tau^{-1}$ defines a morphism of formal schemes. Similarly, to define $\widehat{\tau}^{-1}:\widehat{W}_{s}\to\widehat{\mathcal{V}}_{s}$ it suffices to define a continuous map on completed local rings. For $j=0$ set
$X_{0}=\frac{1}{p\lambda}\bigl(\exp(p\lambda X_{0}^{\prime})-1\bigr)=\sum_{n\geq 1}\frac{(p\lambda)^{n-1}}{n!}(X_{0}^{\prime})^{n}.$
As $p\lambda\in pR$, one has $\nu((p\lambda)^{n-1})\geq n-1\geq\nu(n!)$, hence $X_{0}\in R\llbracket X_{0}^{\prime}\rrbracket$ and $X_{0}\in I_{W}$.
Assume inductively that $X_{0},\ldots,X_{j-1}\in R\llbracket X_{0}^{\prime},\ldots,X_{j-1}^{\prime}\rrbracket$ lie in $I_{W}$, and that after substituting these series into the forward formulas, one recovers $X_{0}^{\prime},\ldots,X_{j-1}^{\prime}$. Define $X_{j}$ by the recursion (40):
$X_{j}=\frac{1}{p\lambda}\Bigl(\exp(p\lambda X_{j}^{\prime})\,G_{j+1,p}(X_{0}^{\prime},\ldots,X_{j-1}^{\prime},0)-G_{j+1}(X_{0},\ldots,X_{j-1})\Bigr).$
Write the numerator as
$(\exp(p\lambda X_{j}^{\prime})-1)\,G_{j+1,p}(X_{0}^{\prime},\ldots,X_{j-1}^{\prime},0)\;+\;\Bigl(G_{j+1,p}(X_{0}^{\prime},\ldots,X_{j-1}^{\prime},0)-G_{j+1}(X_{0},\ldots,X_{j-1})\Bigr).$
The first term is in $(p\lambda)$. For the second term, recall from Proposition 3.8 that
$G^{\prime}_{j+1}(X):=G_{j+1,p}(X_{0}^{\prime}(X),\ldots,X_{j-1}^{\prime}(X),0)-G_{j+1}(X_{0},\ldots,X_{j-1})\in(p\lambda)\,R\llbracket X_{0},\ldots,X_{j-1}\rrbracket.$
Substituting $X_{i}=X_{i}(X_{0}^{\prime},\ldots,X_{i}^{\prime})$ and using the inductive identities $X_{i}^{\prime}(X_{0},\ldots,X_{i})=X_{i}^{\prime}$ for $i<j$, we get
$G_{j+1,p}(X_{0}^{\prime},\ldots,X_{j-1}^{\prime},0)-G_{j+1}(X_{0},\ldots,X_{j-1})\in(p\lambda).$
Hence the numerator lies in $(p\lambda)$, so $X_{j}\in R\llbracket X_{0}^{\prime},\ldots,X_{j}^{\prime}\rrbracket$. Since both terms have constant term $1$, their difference has constant term 0, so $X_{j}\in I_{W}$.
Step 3: the maps are inverse. We check that $\widehat{\tau}^{\sharp}$ and $(\widehat{\tau}^{-1})^{\sharp}$ are mutually inverse by verifying on the topological generators. Let $I_{\mathcal{V}}:=(X_{0},\ldots,X_{s-1})\subset R\llbracket X_{0},\ldots,X_{s-1}\rrbracket$ and $I_{W}:=(X_{0}^{\prime},\ldots,X_{s-1}^{\prime})\subset R\llbracket X_{0}^{\prime},\ldots,X_{s-1}^{\prime}\rrbracket$.
The identities
$\exp(\ln(1+T))=1+T,\qquad\ln(\exp(T))=T,$
hold as formal identities in $R\llbracket T\rrbracket$; hence they hold after substitution for $T\in I_{\mathcal{V}}$ (resp. $T\in I_{W}$) in the $I_{\mathcal{V}}$-adic (resp. $I_{W}$-adic) topology.
For $j=0$, we have
$X_{0}^{\prime}=\frac{1}{p\lambda}\ln(1+p\lambda X_{0}),\qquad X_{0}=\frac{1}{p\lambda}(\exp(p\lambda X_{0}^{\prime})-1),$
so composing in either order reduces immediately to the identities above with $T=p\lambda X_{0}$ (resp. $T=p\lambda X_{0}^{\prime}$).
Assume inductively that the two compositions are the identity on $X^{\prime}_{0},\ldots,X^{\prime}_{j-1}$ (equivalently on $X_{0},\ldots,X_{j-1}$). Set
$\Theta_{j}\,:=\,G_{j+1,p}(X^{\prime}_{0},\ldots,X^{\prime}_{j-1},0)\in R\llbracket X^{\prime}_{0},\ldots,X^{\prime}_{j-1}\rrbracket,\qquad\widetilde{\Theta}_{j}\,:=\,\widehat{\tau}^{\sharp}(\Theta_{j})\in R\llbracket X_{0},\ldots,X_{j-1}\rrbracket.$
Both $\Theta_{j}$ and $\widetilde{\Theta}_{j}$ have constant term $1$, hence are units.
(1) Check $(\widehat{\tau}^{-1})^{\sharp}\circ\widehat{\tau}^{\sharp}(X^{\prime}_{j})=X^{\prime}_{j}$. Set
$V_{j}\,:=\,\frac{G_{j+1}(X_{0},\ldots,X_{j-1})+p\lambda X_{j}}{\widetilde{\Theta}_{j}}\in R\llbracket X_{0},\ldots,X_{j}\rrbracket,$
so by definition
$\widehat{\tau}^{\sharp}(X^{\prime}_{j})=\frac{1}{p\lambda}\ln(V_{j}).$
By Proposition 3.8, $V_{j}=1+p\lambda\Delta_{j}$, so the logarithm is defined. Applying $(\widehat{\tau}^{-1})^{\sharp}$ and using the induction hypothesis on lower coordinates gives
$(\widehat{\tau}^{-1})^{\sharp}\!\bigl(G_{j+1}(X_{0},\ldots,X_{j-1})\bigr)=G_{j+1}\bigl((\widehat{\tau}^{-1})^{\sharp}(X_{0}),\ldots,(\widehat{\tau}^{-1})^{\sharp}(X_{j-1})\bigr),$
and similarly $(\widehat{\tau}^{-1})^{\sharp}(\widetilde{\Theta}_{j})=\Theta_{j}$. By the recursion (40),
$p\lambda\,(\widehat{\tau}^{-1})^{\sharp}(X_{j})=\exp(p\lambda X^{\prime}_{j})\,\Theta_{j}-G_{j+1}\bigl((\widehat{\tau}^{-1})^{\sharp}(X_{0}),\ldots,(\widehat{\tau}^{-1})^{\sharp}(X_{j-1})\bigr).$
Substituting this into $(\widehat{\tau}^{-1})^{\sharp}(V_{j})$ yields
$(\widehat{\tau}^{-1})^{\sharp}(V_{j})=\frac{G_{j+1}(\cdots)+p\lambda\,(\widehat{\tau}^{-1})^{\sharp}(X_{j})}{\Theta_{j}}=\frac{G_{j+1}(\cdots)+\bigl(\exp(p\lambda X^{\prime}_{j})\,\Theta_{j}-G_{j+1}(\cdots)\bigr)}{\Theta_{j}}=\exp(p\lambda X^{\prime}_{j}).$
Therefore
$(\widehat{\tau}^{-1})^{\sharp}\bigl(\widehat{\tau}^{\sharp}(X^{\prime}_{j})\bigr)=\frac{1}{p\lambda}\ln\!\bigl((\widehat{\tau}^{-1})^{\sharp}(V_{j})\bigr)=\frac{1}{p\lambda}\ln\!\bigl(\exp(p\lambda X^{\prime}_{j})\bigr)=X^{\prime}_{j},$
using $\ln(\exp(T))=T$ for $T=p\lambda X^{\prime}_{j}$ (with the stated valuation condition).
(2) Check $\widehat{\tau}^{\sharp}\circ(\widehat{\tau}^{-1})^{\sharp}(X_{j})=X_{j}$. Start from the recursion (40), viewed as the definition of $(\widehat{\tau}^{-1})^{\sharp}(X_{j})$:
$p\lambda\,(\widehat{\tau}^{-1})^{\sharp}(X_{j})=\exp(p\lambda X^{\prime}_{j})\,\Theta_{j}-G_{j+1}\bigl((\widehat{\tau}^{-1})^{\sharp}(X_{0}),\ldots,(\widehat{\tau}^{-1})^{\sharp}(X_{j-1})\bigr).$
Apply $\widehat{\tau}^{\sharp}$ to both sides. By the induction hypothesis, $\widehat{\tau}^{\sharp}\circ(\widehat{\tau}^{-1})^{\sharp}$ fixes $X_{0},\ldots,X_{j-1}$, hence it fixes the polynomial $G_{j+1}(X_{0},\ldots,X_{j-1})$. Also $\widehat{\tau}^{\sharp}(\Theta_{j})=\widetilde{\Theta}_{j}$, and $\widehat{\tau}^{\sharp}(\exp(p\lambda X^{\prime}_{j}))=\exp\bigl(p\lambda\,\widehat{\tau}^{\sharp}(X^{\prime}_{j})\bigr)$. But $\widehat{\tau}^{\sharp}(X^{\prime}_{j})$ was defined so that
$\exp\bigl(p\lambda\,\widehat{\tau}^{\sharp}(X^{\prime}_{j})\bigr)=V_{j}.$
Since $\widehat{\tau}^{\sharp}(\Theta_{j})=\widetilde{\Theta}_{j}$, we get
$\widehat{\tau}^{\sharp}\bigl(\exp(p\lambda X^{\prime}_{j})\,\Theta_{j}\bigr)=\exp\bigl(p\lambda\,\widehat{\tau}^{\sharp}(X^{\prime}_{j})\bigr)\,\widetilde{\Theta}_{j}=V_{j}\,\widetilde{\Theta}_{j}.$
Therefore
$p\lambda\cdot\widehat{\tau}^{\sharp}\bigl((\widehat{\tau}^{-1})^{\sharp}(X_{j})\bigr)=V_{j}\,\widetilde{\Theta}_{j}-G_{j+1}(X_{0},\ldots,X_{j-1})=p\lambda X_{j},$
so dividing by $p\lambda$ gives $\widehat{\tau}^{\sharp}\circ(\widehat{\tau}^{-1})^{\sharp}(X_{j})=X_{j}$. This completes the induction for $j$.
Thus $\widehat{\tau}$ is an isomorphism of formal schemes with inverse induced by (40). ∎
@@
We can therefore naturally declare the group operations on the formal completions $\widehat{\mathcal{W}}_{s}$ and $\widehat{\mathcal{V}}_{s}$ over $R$.
@@
Define a group structure on $\widehat{\mathcal{W}}_{s}$ (and similarly on $\widehat{\mathcal{V}}_{s}$) over $R$ by pulling back the Witt vector group structure on $\widehat{W}_{s}$ via the coordinate maps $\gamma$ (or $\tau$).
@@
The map $\tau$ (resp. $\gamma$) induces a formal group scheme structure over $R$ on the formal completion $\widehat{\mathcal{V}}_{s}$ (resp. $\widehat{\mathcal{W}}_{s}$) along the identity section. Equivalently, the structure maps
$\widehat{\mathcal{V}}_{s}\times_{R}\widehat{\mathcal{V}}_{s}\longrightarrow\widehat{\mathcal{V}}_{s},\qquad\widehat{\mathcal{V}}_{s}\longrightarrow\widehat{\mathcal{V}}_{s},\qquad\operatorname{Spf}(R)\longrightarrow\widehat{\mathcal{V}}_{s}$
(and similarly for $\widehat{\mathcal{W}}_{s}$) are given in coordinates by $s$-tuples of formal power series with coefficients in $R$.